OpenStudy (mathlegend):

Solve the Equation:

OpenStudy (mathlegend):

x/2 + 1/6 = x/3 (ALL FRACTIONS)

OpenStudy (mathlegend):

I wonder could I multiply one of the denominators?

OpenStudy (anonymous):

You can multiply the entire equation by the lowest common denominator, that will eliminate the fractions.

OpenStudy (mathlegend):

by 6?

OpenStudy (anonymous):

Great thinking!!

OpenStudy (anonymous):

Yes.

OpenStudy (mathlegend):

But how would that elminate the x/2?

OpenStudy (anonymous):

Can 6 be divided by 2?

OpenStudy (mathlegend):

Yeah but we are multiplying them right?

OpenStudy (mathlegend):

6 * 2?

OpenStudy (anonymous):

Look at it this way: \(\dfrac{6}{1} \times \dfrac{x}{2}\) You can cancel across the multiplication sign.

OpenStudy (mathlegend):

so I'm multiplying everything by the reciprocal of 1/6

OpenStudy (anonymous):

I think I may have confused you. Let's see if Mertsj can clear it up....

OpenStudy (mertsj):